# A General Control-Theoretic Approach for Reinforcement Learning: Theory and Algorithms

### v1

@@ 2 Control-Based Reinforcement Learning Approach
Recall that the stationary control policy ${\mathcal{F}}_{p}:{{\mathbb{X}}}\rightarrow{\mathbb{W}}$ defines a probability distribution over all available control actions $u\in{\mathbb{U}}$ given the current state $x\in{{\mathbb{X}}}$. Define ${\mathcal{F}}_{p,u}(x)$ to be the element of ${\mathcal{F}}_{p}(x)$ corresponding to action $u$ (i.e., ${\mathcal{F}}_{p,u}(x):=\mathbb{P}[u_{t}=u\,|\,x_{t}=x,{\mathcal{F}}_{p}]$), and correspondingly define $\tilde{q}_{u}(x,{\mathcal{F}}_{p}):=q_{u}(x,{\mathcal{F}}_{p}(x)):=q(x,u)$. Let $x_{0}$ denote the start state at time $t=0$ and $\mathbb{P}_{{\mathcal{F}}_{p}}(x_{0},x,k)$ the probability of going from state $x_{0}$ to state $x$ in $k$ steps under the control policy ${\mathcal{F}}_{p}$. Our general control-policy-parameter gradient ascent result is then formally expressed as follows.
Theorem 2.2.
@@
Consider a family of control policy functions ${\mathbb{F}}$ and its parameter set ${{\mathbb{S}}}$ with contraction operator ${{\mathbf{T}}}$ in the form of (2). We then have
@@
$\nabla_{p}V_{{{\mathcal{F}}_{p}}}(x_{0})=\sum_{x\in{{\mathbb{X}}}}\>\sum_{k=0}^{\infty}\gamma^{k}\left(\mathbb{P}_{{\mathcal{F}}_{p}}(x_{0},x,k)\right)\>\sum_{u\in{\mathbb{U}}}\frac{\partial{\mathcal{F}}_{p,u}(x)}{\partial{\mathcal{F}}_{p}}\frac{\partial{\mathcal{F}}_{p}(x)}{\partial p}\tilde{q}_{u}(x,{{\mathcal{F}}}_{p}),$ (4)
where it is assumed that ${\mathcal{F}}_{p,u}(x)$ is differentiable w.r.t. ${\mathcal{F}}_{p}$ and ${\mathcal{F}}_{p}(x)$ is differentiable w.r.t. $p$, i.e., both $\partial{\mathcal{F}}_{p,u}(x)/\partial{\mathcal{F}}_{p}$ and $\partial{\mathcal{F}}_{p}(x)/\partial p$ exist.
@@
The corresponding gradient ascent result for the special case of deterministic control policies follows directly from Theorem 2.2 with the conditional probability distribution ${\mathcal{F}}_{p,u}(x)$ given by the associated indicator function ${\mathbb{I}}_{{\mathcal{F}}_{p}(x)}(u)$, returning $1$ if and only if $u={\mathcal{F}}_{p}(x)$ and zero otherwise.
@@
Following along similar lines to those of various forms of policy-gradient ascent methods based on the standard policy gradient theorem, we devise control-policy-parameter gradient ascent methods within the context of our CBRL approach based on Theorem 2.2. One such gradient ascent method comprises an iterative process for directly learning the unknown vector $p$ of the optimal control policy w.r.t. the value function $V$ whose iterations, with stepsize $\eta$, proceed according to
$p_{t+1}=p_{t}+\eta\nabla_{p}V_{{{\mathcal{F}}_{p}}}(x_{t}),$ (5)
@@
where $\nabla_{p}V_{{{\mathcal{F}}_{p}}}$ is as given by (4). We note that the standard policy gradient method is a special case of (5) where the parameter vector $p$ is directly replaced by the policy $\pi$ (see, e.g., [21]).
3 Adaptation of Control-Based Reinforcement Learning Approach
@@
Consider our CBRL approach within the context of the LQR control-theoretic framework where at time $t$ the system is in state $x_{t}$ and the parameter vector estimate is $p_{t}$. The optimal (deterministic) control policy is then given by ${\mathcal{F}}_{p}(x_{t})=-Kx_{t}$ via the corresponding solution of the CARE (6), together with the change of coordinates $\tilde{x}=x-x^{*}$ in general when the target $x^{*}$ is not the origin. We then update the parameter vector according to the iterative process (5) where $\nabla_{p}V_{{{\mathcal{F}}_{p}}}$ is obtained from (4) of Theorem 2.2. In particular, the first partial derivative term on the right hand side of (4) is given by the standard policy gradient solution, and the second partial derivative term on the right hand side of (4) is obtained by differentiation of the CARE (6), which in turn renders [27]
@@
Our CBRL approach under the LQR control-theoretic framework concerns the linear dynamics $\dot{x}=Ax+Bu$ where $A$ and $B$ contain elements of the unknown parameter vector $p$. By leveraging known basic information about the LQR control problem at hand, only a relatively small number of unknown parameters need to be learned. For robotics and physics applications where state variables are derivatives of each other w.r.t. time (e.g., position, velocity, acceleration, jerk), the corresponding rows in the $A$ matrix consist of a single $1$ and the corresponding rows in $B$ comprise zeros. We exploit this basic information to consider general matrix forms for $A$ and $B$ that reduce the number of unknown parameters to be learned. As a representative illustration, the system dynamics when there are 2 groups of such variables have the form given in (9) of Appendix B.
4 Experimental Results
@@
In this section, we conduct numerical experiments to evaluate the performance of our CBRL approach of Section 2 adapted to the LQR control-theoretic framework according to Section 3, as summarized in Algorithm 1 of Appendix A.4. Our numerical experiments consider several classical control tasks from Gymnasium [28], including Cart Pole, Lunar Lander (Continuous), Mountain Car (Continuous), and Pendulum. We compare our CBRL method with three state-of-the-art RL algorithms, namely DQN [7] for discrete actions, DDPG [29] for continuous actions and PPO [23], together with a variant of PPO that solely replaces the nonlinear policy of PPO with a linear policy (since we know the optimal policy for some of the problems, e.g., Cart Pole, is linear). These baselines are selected as the state-of-the-art RL algorithms for solving the control tasks under consideration. Experimental details and additional results, including validation of the robustness of our CBRL approach, are provided in Appendix C both in general and for each control task.
@@
Our CBRL approach depends upon the control-theoretic framework to which it is adapted, where we have chosen LQR as a representative example with the understanding that not all control tasks can be adequately solved using LQR even if the parameter vector $p$ is known. Recall, however, that our CBRL approach allows the domain of the control policies to span a subset of states in ${{\mathbb{X}}}$, thus enabling partitioning of the state space so that properly increased richness w.r.t. finer and finer granularity can provide improved approximations and asymptotic optimality according to Theorem 2.1, analogous to canonical piecewise-linear approximations [25]. While Cart Pole and Lunar Lander (Continuous) can be directly addressed within the context of LQR, this is not the case for Mountain Car (Continuous) and Pendulum which require a nonlinear controller. We therefore partition the state space in the case of such control tasks and consider a corresponding piecewise-LQR controller where the learned parameter vectors may differ or be shared across the partitions.
(a) Return v.s. Episode
(b) Return v.s. Time
@@
Figure 1: Learning curves of CartPole-v0 over five independent runs. (a) and (b): Return vs. number of episodes and running time, respectively, for our CBRL approach in comparison with the Linear policy, PPO, and DQN. (c): Learning behavior of CBRL parameters.
(a) Return v.s. Episode
(b) Return v.s. Time
@@
Figure 2: Learning curves of LunarLanderContinuous-v2 over five independent runs. (a) and (b): Return vs. number of episodes and running time, respectively, for our CBRL approach in comparison with the Linear policy, PPO, and DDPG. (c): Learning behavior of CBRL parameters.
(a) Return v.s. Episode
(b) Return v.s. Time
@@
Figure 3: Learning curves of MountainCarContinuous-v0 over five independent runs. (a) and (b): Return vs. number of episodes and running time, respectively, for our CBRL approach in comparison with the Linear policy, PPO, and DDPG. (c): Learning behavior of CBRL parameters.
(a) Return v.s. Episode
(b) Return v.s. Time
@@
Figure 4: Learning curves of Pendulum-v1 over five independent runs. (a) and (b): Return vs. number of episodes and running time, respectively, for our CBRL approach in comparison with the Linear policy, PPO, and DDPG. (c): Learning behavior of CBRL parameters.
4.1 Cart Pole under CBRL LQR
Cart Pole consists of a pole connected to a horizontally moving cart with the goal of balancing the pole by applying a force on the cart to the left or right. The state of the system is $x=[s,\dot{s},\theta,\dot{\theta}]$ in terms of the position of the cart $s$, the velocity of the cart $\dot{s}$, the angle of the pole $\theta$, and the angular velocity of the pole $\dot{\theta}$. With upright initialization of the pole on the cart, each episode comprises $200$ steps and the problem is considered “solved” upon achieving an average return of $195$.
The LQR controller can be used to solve the Cart Pole problem provided that all the parameters of the system are known (e.g., mass of cart, mass and length of pole, gravity). We address the problem within the context of our CBRL approach by exploiting the general matrix form for the LQR dynamics (10) (Appendix C.1), solely taking into account general physical relationships and laws, with the unknown parameters $a_{0},\ldots,a_{7},b_{0},b_{1}$ to be learned (Fig. 1(c)).
Fig. 1 presents numerical results for the three state-of-the-art baselines (discrete actions) and our CBRL approach, with each run over five independent random seeds. All parameters in (10) are initialized uniformly in $(0,1)$ and then learned using our policy-parameter gradient ascent iteration (5). Fig. 1(a)–(b) clearly demonstrate that our CBRL approach outperforms all other baselines w.r.t. both the number of episodes and running time, as well as demonstrating a more stable training process.
4.2 Lunar Lander (Continuous) under CBRL LQR
Lunar Lander (Continuous) is a classical spacecraft trajectory optimization problem with the goal to land softly and fuel-efficiently on a landing pad by applying thrusters to the left, to the right and upward. The state of the system is $x=[s_{x},v_{x},s_{y},v_{y},\theta,\dot{\theta}]$ in terms of $(x,y)$ positions $s_{x}$ and $s_{y}$, two linear velocities $v_{x}$ and $v_{y}$, angle $\theta$, and angular velocity $\dot{\theta}$. Starting at the top center of the viewport with a random initial force applied to its center of mass, the lander (spacecraft) is subject to gravity, friction and turbulence while surfaces on the “moon” are randomly generated in each episode with the target landing pad centered at $(0,0)$. The problem is considered “solved” upon achieving an average return of $200$.
The LQR controller can be used to solve the Lunar Lander (Continuous) problem provided that all the parameters of the system are known (e.g., mass of lander, gravity, friction, etc). We address the problem within the context of our CBRL approach by exploiting the general matrix form for the LQR dynamics (11) (Appendix C.2), solely taking into account mild physical information from the system state, with the unknown parameters $a_{0},\ldots,a_{11},b_{0},b_{1},b_{2}$ to be learned (Fig. 2(c)).
Fig. 2 presents numerical results for the three state-of-the-art baselines (continuous actions) and our CBRL approach, with each run over five independent random seeds. All parameters in (11) are initialized uniformly within $(0,1)$ and then learned using our control-policy-parameter gradient ascent iteration (5). Fig. 2(a)–(b) clearly demonstrate that our CBRL approach outperforms all other baselines w.r.t. both the number of episodes and running time, in addition to demonstrating a more stable training process. We note that the baseline algorithms often crash and terminate sooner than the more successful landings of our CBRL approach, resulting in the significantly worse performance exhibited in Fig. 2(a)–(b).
4.3 Mountain Car (Continuous) under CBRL Piecewise-LQR
@@
Mountain Car (Continuous) consists of the placement of a car in a valley with the goal of accelerating the car to reach the target at the top of the hill on the right by applying a force on the car to the left or right. The state of the system is $x=[s,v]$ in terms of the position of the car $s$ and the velocity of the car $v$. With random initialization at the bottom of the valley, the problem is considered “solved” upon achieving an average return of $90$.
@@
Recalling that the LQR controller is not sufficient to solve this problem, we consider a piecewise-LQR controller that partitions the state space into two regions (see Fig. 9 in Appendix C.3). The target state $x^{*}=[s^{*},v^{*}]$ is selected to be $x^{*}=[-1.2,0]$ if $v<0$, and to be $x^{*}=[0.6,0]$ otherwise, where $-1.2$ and $0.6$ represent the position of the left hill and the right hill, respectively. We address the problem within the context of our CBRL approach by exploiting the general matrix form for the piecewise-LQR dynamics (12) (Appendix C.3), solely taking into account general physical relationships and laws, with the unknown parameters $a_{0},a_{1},b_{0},c_{0}$ to be learned (Fig. 3(c)).
@@
Fig. 3 presents numerical results for the three state-of-the-art (continuous actions) baselines and our CBRL approach, with each run over five independent random seeds. All parameters in (12) are initialized uniformly within $(0,1)$ and then learned using our control-policy-parameter gradient ascent iteration (5). Fig. 3(a)–(b) clearly demonstrate that our CBRL approach outperforms all other baselines w.r.t. both the number of episodes and running time, in addition to demonstrating a more stable training process.
4.4 Pendulum under CBRL Piecewise-LQR
@@
Pendulum consists of a link attached at one end to a fixed point and the other end being free, with the goal of swinging up to an upright position by applying a torque on the free end. The state of the system is $x=[\theta,\dot{\theta}]$ in terms of the angle of the link $\theta$ and the angular velocity of the link $\dot{\theta}$. With random initial position of the link, each episode comprises $200$ steps and the problem is considered “solved” upon achieving an average return of $-200$.

### v2

@@ Introduction
Recall that the stationary control policy ${\mathcal{F}}_{p}:{{\mathbb{X}}}\rightarrow{\mathbb{W}}$ defines a probability distribution over all available control actions $u\in{\mathbb{U}}$ given the current state $x\in{{\mathbb{X}}}$. Define ${\mathcal{F}}_{p,u}(x)$ to be the element of ${\mathcal{F}}_{p}(x)$ corresponding to action $u$ (i.e., ${\mathcal{F}}_{p,u}(x):=\mathbb{P}[u_{t}=u\,|\,x_{t}=x,{\mathcal{F}}_{p}]$), and correspondingly define $\tilde{q}_{u}(x,{\mathcal{F}}_{p}):=q_{u}(x,{\mathcal{F}}_{p}(x)):=q^{p}(x,u)$, where the latter denotes the $Q$-function of the policy ${\mathcal{F}}_{p}$. Let $x_{0}$ denote the start state at time $t=0$ and $\mathbb{P}(x,k\,|\,x_{0},{\mathcal{F}}_{p})$ the probability of going from state $x_{0}$ to state $x$ in $k$ steps under the control policy ${\mathcal{F}}_{p}$. Our general control-policy-parameter gradient ascent result is then formally expressed as follows.
Theorem 3.
@@
Consider a family of control policy functions ${\mathbb{F}}$, its parameter set ${{\mathbb{S}}}$ with contraction operator ${{\mathbf{T}}}$ in the form of (CBRL Approach), and the value function $V_{{{\mathcal{F}}_{p}}}$ under the control policy ${\mathcal{F}}_{p}$. Assuming ${\mathcal{F}}_{p,u}(x)$ is differentiable w.r.t. ${\mathcal{F}}_{p}$ and ${\mathcal{F}}_{p}(x)$ is differentiable w.r.t. $p$, we then have
$\nabla_{p}V_{{{\mathcal{F}}_{p}}}(x_{0})=$ $\sum_{x\in{{\mathbb{X}}}}\>\sum_{k=0}^{\infty}\gamma^{k}\,\mathbb{P}(x,k\,|\,x_{0},{\mathcal{F}}_{p})\;\;\times$
@@
$\quad\left[\sum_{u\in{\mathbb{U}}}\frac{\partial{\mathcal{F}}_{p,u}(x)}{\partial{\mathcal{F}}_{p}}\frac{\partial{\mathcal{F}}_{p}(x)}{\partial p}\tilde{q}_{u}(x,{{\mathcal{F}}}_{p})\right].$ (5)
@@
The corresponding gradient ascent result for the case of deterministic control policies follows directly from Theorem 3 with the conditional probability distribution ${\mathcal{F}}_{p,u}(x)$ given by the associated indicator function ${\mathbb{I}}_{{\mathcal{F}}_{p}(x)}(u)$, returning $1$ if and only if $u={\mathcal{F}}_{p}(x)$ and zero otherwise.
@@
Following along similar lines of the various forms of policy-gradient ascent methods based on the standard policy gradient theorem, we devise control-policy-parameter gradient ascent methods within the context of our CBRL approach based on Theorem 3. One such gradient ascent method comprises an iterative process for directly learning the unknown parameter vector $p$ of the optimal control policy w.r.t. the value function $V_{{\mathcal{F}}_{p}}$ whose iterations, with stepsize $\eta$ and $\nabla_{p}V_{{{\mathcal{F}}_{p}}}$ given by (3), proceed according to
$p_{t+1}=p_{t}+\eta\nabla_{p}V_{{{\mathcal{F}}_{p}}}(x_{t}).$ (6)
@@
Note that standard policy gradient methods are special cases of (6) where the parameter vector $p_{t}$ is directly replaced by the policy $\pi_{t}$ (see, e.g., (Agarwal et al. 2021)).
Adaptation of CBRL Approach
@@
Consider our CBRL approach within the context of the LQR control-theoretic framework where at time $t$ the system is in state $x_{t}$ and the parameter vector estimate is $p_{t}$. The optimal (deterministic) control policy is then given by ${\mathcal{F}}_{p}(x_{t})=-Kx_{t}$ via the corresponding solution of the CARE (7), together with the change of coordinates $\tilde{x}=x-x^{*}$ in general when the target $x^{*}$ is not the origin. We then update the parameter vector according to the iterative process (6) where $\nabla_{p}V_{{{\mathcal{F}}_{p}}}$ is obtained from (3) of Theorem 3. In particular, the first partial derivative term on the right hand side of (3) is given by the standard policy gradient solution, and the second partial derivative term on the right hand side of (3) is obtained by differentiation of the CARE (7), which in turn renders (Kao and Hennequin 2020)
@@
Our CBRL approach under the LQR control-theoretic framework concerns the linear dynamics $\dot{x}=Ax+Bu$ where $A$ and $B$ contain elements of the unknown parameter vector $p$. By leveraging known basic information about the LQR control problem at hand, only a relatively small number of unknown parameters need to be learned. For a wide range of applications where state variables are derivatives of each other w.r.t. time (e.g., position, velocity, acceleration, jerk), the corresponding rows in the $A$ matrix consist of a single $1$ and the corresponding rows in $B$ comprise zeros. We exploit this basic information to consider general matrix forms for $A$ and $B$ that reduce the number of unknown parameters to be learned. As a representative illustration, the system dynamics when there are two groups of such variables have the form given by
$\underbrace{\begin{bmatrix}\frac{dx_{1}}{dt}\\ \vdots\\ \frac{d^{k_{1}}x_{1}}{dt^{k_{1}}}\\ \frac{dx_{2}}{dt}\\ \vdots\\ \frac{d^{k_{2}}x_{2}}{dt^{k_{2}}}\end{bmatrix}}_{\dot{x}}\!\!=\!\!\underbrace{\begin{bmatrix}0&1&0&\cdots&&&0\\ 0&0&1&\cdots&&\\ a_{10}&\cdots&&&&&a_{1n}\\ 0&\cdots&&0&1&\dots&0\\ 0&\cdots&&&0&1&\cdots\\ a_{20}&\cdots&&&&&a_{2n}\end{bmatrix}}_{A}x+\underbrace{\begin{bmatrix}0\\ \vdots\\ b_{0}\\ 0\\ \vdots\\ b_{1}\end{bmatrix}}_{B}u.$
Experimental Results
@@
In this section, we conduct numerical experiments to evaluate the performance of our CBRL approach adapted to the above LQR control-theoretic framework, as summarized in Algorithm 1 of the appendix. The objective (1) seeks to maximize the expected cumulative discounted reward, or return. Our numerical experiments consider several classical RL tasks from Gymnasium (Towers et al. 2023), including Cart Pole, Lunar Lander (Continuous), Mountain Car (Continuous), and Pendulum. We compare our CBRL method with three state-of-the-art RL algorithms, namely DQN (Mnih et al. 2013) for discrete actions, DDPG (Lillicrap et al. 2015) for continuous actions and PPO (Schulman et al. 2017), together with a variant of PPO that solely replaces the nonlinear policy of PPO with a linear policy (since we know the optimal policy for some of the problems, e.g., Cart Pole, is linear). These baselines are selected as the state-of-the-art algorithms for solving the RL tasks under consideration. Experimental details and additional results are provided in the appendix both in general and for each RL task.
@@
Our CBRL approach depends upon the control-theoretic framework to which it is adapted, where we have chosen LQR as a representative example with the understanding that not all RL tasks can be adequately solved using LQR even if the parameter vector $p$ is known. Recall, however, that our CBRL approach allows the domain of the control policies to span a subset of states in ${{\mathbb{X}}}$, thus enabling partitioning of the state space so that properly increased richness w.r.t. finer and finer granularity can provide improved approximations and asymptotic optimality according to Theorem 2, analogous to the class of canonical piecewise-linear approximations (Lin and Unbehauen 1992). While Cart Pole and Lunar Lander can be directly addressed within the context of LQR, this is not the case for Mountain Car and Pendulum which require a nonlinear controller. We therefore partition the state space in the case of such RL tasks and consider a corresponding piecewise-LQR controller where the learned parameter vectors may differ or be shared across the partitions.
All parameters of our CBRL approach are randomly initialized uniformly within $(0,1)$ and then learned using our control-policy-parameter gradient ascent iteration (6). We consider four sets of initial parameters to validate the robustness of our CBRL approach for each RL task (see appendix).
Cart Pole under CBRL LQR
Cart Pole, as depicted in Fig. 5 of the appendix, consists of a pole connected to a horizontally moving cart with the goal of balancing the pole by applying a force on the cart to the left or right. The state of the system is $x=[s,\dot{s},\theta,\dot{\theta}]$ in terms of the position of the cart $s$, the velocity of the cart $\dot{s}$, the angle of the pole $\theta$, and the angular velocity of the pole $\dot{\theta}$. With upright initialization of the pole on the cart, each episode comprises $200$ steps and the problem is considered “solved” upon achieving an average return of $195$.
The LQR controller can be used to solve the Cart Pole problem provided that all the parameters of the system are known (e.g., mass of cart, mass and length of pole, gravity), where the angle is kept small by our CBRL approach adapted to the LQR framework. We address the problem within the context of our CBRL approach by exploiting the general matrix form for the LQR dynamics given by (15) in the appendix, solely taking into account general physical relationships (e.g., the derivative of the angle of the pole is equivalent to its angular velocity) and laws (e.g., the force can only affect the acceleration), with the unknown parameters $a_{0},\ldots,a_{7},b_{0},b_{1}$ to be learned; refer to Fig. 1(c).
Fig. 1 and Table 1 (appendix) present numerical results for the three state-of-the-art baselines (discrete actions) and our CBRL approach, with each run over five independent random seeds; see Table 2 and Fig. 6 in the appendix for the four sets of initial parameters in (15). Fig. 1(a) – (b), Table 1 and Fig. 6(a) – (b) clearly demonstrate that our CBRL approach provides far superior performance (both mean and std. dev.) over all baselines w.r.t. both the number of episodes and running time, in addition to demonstrating a more stable training process. Fig. 6(c) – (f) illustrates the learning behavior of CBRL parameters given different initialization.
Lunar Lander under CBRL LQR
Lunar Lander (Continuous), as depicted in Fig. 7 of the appendix, is a classical spacecraft trajectory optimization problem with the goal to land softly and fuel-efficiently on a landing pad by applying thrusters to the left, to the right, and upward. The state of the system is $x=[s_{x},v_{x},s_{y},v_{y},\theta,\dot{\theta}]$ in terms of the $(x,y)$ positions $s_{x}$ and $s_{y}$, two linear velocities $v_{x}$ and $v_{y}$, angle $\theta$, and angular velocity $\dot{\theta}$. Starting at the top center of the viewport with a random initial force applied to its center of mass, the lander (spacecraft) is subject to gravity, friction and turbulence while surfaces on the “moon” are randomly generated in each episode with the target landing pad centered at $(0,0)$. The problem is considered “solved” upon achieving an average return of $200$.
The LQR controller can be used to solve the Lunar Lander (Continuous) problem provided that all parameters of the system are known (e.g., mass of lander, gravity, friction, etc.), where the angle is kept small by our CBRL approach adapted to the LQR framework. We address the problem within the context of our CBRL approach by exploiting the general matrix form for the LQR dynamics given by (16) in the appendix, solely taking into account general physical relationships (akin to Cart Pole) and mild physical information from the system state (e.g., the acceleration is independent of the position), with the unknown parameters $a_{0},\ldots,a_{11},b_{0},b_{1},b_{2}$ to be learned; refer to Fig. 2(c).
Fig. 2 and Table 3 (see appendix) present numerical results for the three state-of-the-art baselines (continuous actions) and our CBRL approach, with each run over five independent random seeds; see Table 4 and Fig. 8 in the appendix for the four sets of initial parameters in (16). Fig. 2(a) – (b), Table 3 and Fig. 8(a) – (b) clearly demonstrate that our CBRL approach provides far superior performance (both mean and std. dev.) over all baselines w.r.t. both the number of episodes and running time, in addition to demonstrating a more stable training process. We note that the baseline algorithms often crash and terminate sooner than the more successful landings of our CBRL approach, resulting in the significantly worse performance exhibited in Fig. 2, Table 3 and Fig. 8. Finally, Fig. 8(c) – 8(f) illustrates the learning behavior of CBRL parameters given different initialization.
(a) Return vs. Episode
(b) Return vs. Time
@@
Figure 1: Learning curves of CartPole-v0 over five independent runs, where the solid line shows the mean and the shaded area depicts the standard deviation. (a) and (b): Return vs. number of episodes and running time, respectively, for our CBRL approach in comparison with the Linear policy, PPO, and DQN. (c): Learning behavior of CBRL parameters.
(a) Return vs. Episode
(b) Return vs. Time
@@
Figure 2: Learning curves of LunarLanderContinuous-v2 over five independent runs, where the solid line shows the mean and the shaded area depicts the standard deviation. (a) and (b): Return vs. number of episodes and running time, respectively, for our CBRL approach in comparison with the Linear policy, PPO, and DDPG. (c): Learning behavior of CBRL parameters.
(a) Return vs. Episode
(b) Return vs. Time
@@
Figure 3: Learning curves of MountainCarContinuous-v0 over five independent runs, where the solid line shows the mean and the shaded area depicts the standard deviation. (a) and (b): Return vs. number of episodes and running time, respectively, for our CBRL approach in comparison with the Linear policy, PPO, and DDPG. (c): Learning behavior of CBRL parameters.
(a) Return vs. Episode
(b) Return vs. Time
@@
Figure 4: Learning curves of Pendulum-v1 over five independent runs, where the solid line shows the mean and the shaded area depicts the standard deviation. (a) and (b): Return vs. number of episodes and running time, respectively, for our CBRL approach in comparison with the Linear policy, PPO, and DDPG. (c): Learning behavior of CBRL parameters.
Mountain Car under CBRL Piecewise-LQR
@@
Mountain Car (Continuous), as depicted in Fig. 9 of the appendix, consists of the placement of a car in a valley with the goal of accelerating the car to reach the target at the top of the hill on the right by applying a force on the car to the left or right. The system state is $x=[s,v]$ in terms of the position of the car $s$ and the velocity of the car $v$. With random initialization at the bottom of the valley, the problem is considered “solved” upon achieving an average return of $90$.
@@
Recall that the LQR controller is not sufficient to solve the Mountain Car (Continuous) problem, even if all the parameters of the system are known (e.g., mass of the car and gravity), because a nonlinear controller is required. Consequently, we consider a piecewise-LQR controller that partitions the state space into two regions (LQR 1 and LQR 2 in Fig. 9 of the appendix). The target state $x^{*}=[s^{*},v^{*}]$ is selected to be $x^{*}=[-1.2,0]$ if $v<0$ (LQR 1), and to be $x^{*}=[0.6,0]$ otherwise (LQR 2), where $-1.2$ and $0.6$ represent the position of the left hill and the right hill, respectively. We address the problem within the context of our CBRL approach by exploiting the general matrix form for the piecewise-LQR dynamics given by (17) in the appendix, solely taking into account general physical relationships and laws, with the unknown parameters $a_{0},a_{1},b_{0},c_{0}$ to be learned; refer to Fig. 3(c).
@@
Fig. 3 and Table 5 (see appendix) present numerical results for the three state-of-the-art baselines (continuous actions) and our CBRL approach, with each run over five independent random seeds; refer to Table 6 and Fig. 10 in the appendix for the four sets of initial parameters in (17). Fig. 3(a) – (b), Table 5 and Fig. 10(a) – (b) clearly demonstrate that our CBRL approach provides superior performance (both mean and std. dev.) over all baselines w.r.t. both the number of episodes and running time, in addition to demonstrating a more stable training process. Fig. 10(c) – 10(f) illustrates the learning behavior of CBRL parameters given different initialization.
Pendulum under CBRL Piecewise-LQR
@@
Pendulum, as depicted in Fig. 11 of the appendix, consists of a link attached at one end to a fixed point and the other end being free, with the goal of swinging up to an upright position by applying a torque on the free end. The state of the system is $x=[\theta,\dot{\theta}]$ in terms of the angle of the link $\theta$ and the angular velocity of the link $\dot{\theta}$. With random initial link position, each episode comprises $200$ steps and the problem is “solved” upon achieving an average return of $-200$.